2 passengers with a combined mass of 150kg are observed to compress the springs of a 1200kg car 2cm. What is the vibrational period of the car and it's passengers?


1
Expert's answer
2021-02-03T16:15:17-0500
"T=2\pi\sqrt{\frac{m}{k}}\\T=2\pi\sqrt{\frac{mx}{Mg}}\\
T=2\pi\sqrt{\frac{150\cdot0.02}{1200\cdot9.8}}=0.1\ s"